OpenStudy (lopus):

hi everybody ,What do you consider the best book for learning math from the beginning? aritmetic--> algebra -->baldor geometry ---> calculus --->

OpenStudy (skullpatrol):

Do you want books written English?

OpenStudy (anonymous):

nelson and pearson

OpenStudy (lopus):

it doesn't matter!

OpenStudy (lopus):

what for @skullpatrol?

OpenStudy (anonymous):

Elementary Algebra for Schools Hall and Knight 1906

OpenStudy (anonymous):

sample problem resolve in to factors \(a^2-2ax+x^2-4b^2\)

OpenStudy (skullpatrol):

$$a^2-2ax+x^2-4b^2=(a-x)^2-(2b)^2=(a-x+2b)(a-x-2b)$$